OpenStudy (anonymous):

The regular price of an item is $84. The item is on sale with a discount of $12. What is the discount rate? A. 0.143% B. 1.43% C. 14.3% D. 143%

OpenStudy (anonymous):

hey can someone help me

OpenStudy (anonymous):

i will give a medal

OpenStudy (acxbox22):

we can use proportions to solve for the discount rate where x is the discount rate 12/84=x/100 84x=1200 x=??